OpenStudy (anonymous):

two 10cm diameter charged rings face each other 20cm apart. both rings are charged to +40nC. what is the electric field strength...

OpenStudy (anonymous):

a) at the midpoint between 2 rings? should the E just be added together but using r=10cm?

OpenStudy (earthcitizen):

\[E=F/Q\]

OpenStudy (anonymous):

found the answer on cramster step by step but not understanding the r: Electric field due to charged ring E = KQx / [ x ^ 2 + R^2 ] ^ 1.5

OpenStudy (anonymous):

(x^2+R^2)^1.3??

OpenStudy (anonymous):

oops 1.5

OpenStudy (earthcitizen):

is it between to plates ?

OpenStudy (anonymous):

yes

OpenStudy (earthcitizen):

two plates*

OpenStudy (anonymous):

yes

OpenStudy (earthcitizen):

\[\epsilon _{0}\ = ?]

OpenStudy (anonymous):

8.85*10^-12

OpenStudy (anonymous):

k=1/(4pi8.85*10^-12)=8.99*10^9

OpenStudy (earthcitizen):

first convert cm to m

OpenStudy (anonymous):

right

OpenStudy (anonymous):

.1m diamter, .2m apart

OpenStudy (anonymous):

and +40nC=+

OpenStudy (anonymous):

+40*10^-9

OpenStudy (anonymous):

answer says it should be 0 cuz they cancel each other but i thought it would double since both rings are +

OpenStudy (earthcitizen):

both rings are either + /-

OpenStudy (earthcitizen):

calculate the force, r= d/2

OpenStudy (anonymous):

i actually understand now from the set up on cramster except for the part that is raised to 1.5...typo? should be .5

OpenStudy (anonymous):

o charge density for ring=Q/(x^2+r^2)^1.5

OpenStudy (anonymous):

i mean Q/(2pi*r)

OpenStudy (earthcitizen):

r=d/2

OpenStudy (earthcitizen):

d=0.1m/2 = 0.05m

OpenStudy (earthcitizen):

total net E=E1+E2